OpenStudy (espex):

Homogeneous Differential: Determine whether the given function is homogeneous. If so, state the degree of homogeneity. \[cos\frac{x^2}{x+y}\] The book says that this is not homogeneous, but I do not understand why or how to verify this.

OpenStudy (eyust707):

Is there a differential somewhere?

OpenStudy (espex):

No, the only thing required is to determine if this equation is homogeneous.

OpenStudy (amistre64):

i believe theres some book work that says; if f(tx,ty) = t^n f(x,y) then its homogenous, if not; then it aint

OpenStudy (espex):

Can it simply be that the numerator is a power of 2 and the denominator has powers of 1?

OpenStudy (amistre64):

if\[cos\frac{(tx)^2}{tx+ty}=t^n~cos\frac{x^2}{x+y}\]it would be homogenous

OpenStudy (espex):

Yes, but when you work it out you get \[\frac{t^2}{t} cos\frac{x^2}{x+y}\]

OpenStudy (espex):

By that logic you would think it homogeneous, yet the book says it is not.

OpenStudy (amistre64):

can you show the steps that allows you to pull out a t?

OpenStudy (espex):

@amistre64 I used essentially the same method putting them in. \[cos\frac{(tx)^2}{t(x+y)} \rightarrow cos\frac{t^2x^2}{tx+ty} \rightarrow cos\frac{t^2}{t}\frac{x^2}{x+y}\]

OpenStudy (amistre64):

have you seen the series expansion for this by chance?

OpenStudy (espex):

Would this be \[cos\frac{t^2}{t} \times cos\frac{x^2}{x+y}\]? I have not or do not recall the series expansion.

OpenStudy (amistre64):

\[cos(t\frac{x^2}{x+y})\ne t^ncos(\frac{x^2}{x+y})\] or am i missing something?

OpenStudy (espex):

@amistre64 That makes sense, it seems it was I that was missing it. :) Thanks so much.

OpenStudy (amistre64):

the series expansion shows something to the effect of\[1-\frac{t^2x^4}{2y^2}+\frac{t^2x^5}{y^3}...\] which of course cant factor out the t^2 wither

OpenStudy (amistre64):

good luck ;)